OpenStudy (anonymous):

9 squared + the square root of 115 = 14 squared

OpenStudy (akshay_budhkar):

? there seems to be something wrong in the question

OpenStudy (anonymous):

doesit make a right triangle?

OpenStudy (akshay_budhkar):

oh ok hehe you just need to check if that is true that is if LHS = RHS in this case it is true, so they do make a right angeld triangle using pythagoras

OpenStudy (akshay_budhkar):

\[9^2+ (\sqrt{115})^2= 81+115=196= 14^2\]

OpenStudy (anonymous):

so doest it?

OpenStudy (akshay_budhkar):

Yes

OpenStudy (anonymous):

ok now do the square root of 24 + 7 squared = 9 squared

OpenStudy (lgbasallote):

@Akshay Budhkar...you cannot simply square one term in an equation it would result to inequality